# Cloud Storage Integrity at Scale:

### v2

@@ IV-A. Optimal Definition
We observe that the problem of finding an optimal hash tree can be reduced to the problem of finding an optimal prefix tree (or prefix code) in the context of lossless data compression (Moffat, 2019). Prefix codes map a set of symbols onto a set of codewords, with the goal of compression being that codewords are as short as possible to produce a maximally compressed representation of the original data. An example is shown in Figure 3. More formally:
@@
A hash tree constructed as an optimal prefix code is optimal for a fixed access probability distribution.
@@
Let $A=\{a_{1},a_{2},\ldots,a_{n}\}$ be a symbol alphabet and $W=\{w_{1},w_{2},\ldots,w_{n}\}$ be a set of associated symbol weights. Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be a prefix code that represents the set of codewords for symbols in $A$. A prefix code $C$ is said to be optimal if it minimizes the expected codeword length: $\operatorname*{arg\,min\,}_{C}~{}\sum_{i=1}^{n}w_{i}|c_{i}|,c_{i}\in C$. The length of a codeword is the number of bits in a codeword, or equivalently, the number of edges in the path from the root to the symbol leaf in the prefix tree representation of $C$. Huffman coding is a widely-used algorithm to produce optimal prefix codes (Huffman, 1952).
@@
IV-B. Extended Optimal Definition
Now we extend our optimal definition to consider the effects of cache performance. Note that both data blocks and hashes can be cached in main memory. In the compression domain, codeword paths in an (optimal) prefix tree can be parsed in a constant amount of work per edge, giving a total amount of work: $\sum_{i=1}^{n}w_{i}(|c_{i}|\cdot O(1))=O(1)\cdot\sum_{i=1}^{n}w_{i}|c_{i}|$. However computing a hash requires (at least) two hash fetches in a binary hash tree: the node’s two children. If both hashes are present in memory, fetch costs are negligible and the amount of work is similarly optimal: $\sum_{i=1}^{n}f_{i}(|b_{i}|\cdot O(1))=O(1)\cdot\sum_{i=1}^{n}f_{i}|b_{i}|$. Otherwise if they must be fetched from disk, I/O costs are non-negligible: $\sum_{i=1}^{n}f_{i}(|b_{i}|\cdot t(b_{i}))=\sum_{i=1}^{n}f_{i}|b_{i}|\cdot t(b_{i})$, for some function $t(b_{i})$.
@@ IV-B. Extended Optimal Definition
We also see that with an optimal cache ($m=0.0$), the expected total amount of work is exactly optimal. However, it has been empirically observed that as cache size decreases, miss rates increase with a power law (Chow, 1974), and thus as the cache size decreases, expected I/O costs increase with a power law. This implies that the performance of hash trees is very sensitive to cache size. In particular, as cache memory can be financially costly on cloud servers, being able to synergize well with relatively smaller caches (w.r.t. larger disks) is critical to a practical hash tree deployment.
@@
Moreover, a Huffman tree is optimal under a fixed set of weights (cf. access probability distribution). If the symbol sequence (cf. block access sequence) observed while compressing a message (cf. during a program trace) exactly matches the one used to construct the tree, then the tree will be exactly optimal (i.e., provide optimal throughput). However, if the sequence deviates from the one used to construct the tree, the tree will not be exactly optimal. In this case, I/O costs will also increase at an amount proportional to the distance between the access probability distributions.
V. Optimal Tree Oracle
@@
We now introduce a new method to instantiate an optimal hash tree in an offline setting. Our optimal definition provides that, if we have knowledge of a concrete block access sequence (i.e., workload trace), we can instantiate an optimal hash tree from the trace. In an offline setting, where we have access to workload traces (e.g., recorded with standard tools like blktrace), we can feasibly do so. Replaying the trace against the optimal tree will then determine the maximum performance achievable under the given workload. Replaying the trace against other hash tree designs will reveal how they perform with respect to the optimal.
@@
We refer to this methodology as the optimal tree oracle, and it allows us to measure a concrete upper bound on performance. The primary purpose is to understand whether overheads over the baseline should be attributed to a fundamental scaling limit or to the structure of the tree itself. For example, a hash tree may be performing optimally, but have high overheads, which would suggest that complimentary optimizations (e.g., parallelization) may be the only way to break the performance ceiling. In contrast, a hash tree that does not perform optimally under a given workload may require a fundamental redesign.
@@
We liken this approach to Belady’s optimal page replacement algorithm (Belady, 1966), a clairvoyant algorithm that has a priori knowledge of future memory accesses and can make optimal page replacement decisions. This gives us the ability to make rigorously grounded conclusions about when a hash tree design performs well and when.
Figure 4. We rigorously evaluate performance by using an optimal tree oracle, which assumes knowledge of a workload trace and uses it to construct an optimal hash tree offline.
V-A. Workflow
The above procedure is composed of five steps (Figure 4).
(1) Workload Generation:
Workload traces are first collected from applications or generated with standard benchmark tools like fio. The traces are composed of tuples of the form: $(op,offset,length)$, where $op$ is either read or write.
(2) Parameter Configuration:
The system is then provisioned with the appropriate storage capacity, hash cache size, and read/write-ratio for workloads. The hash cache size is configured as a percentage of the total number of nodes in the hash tree. For example, a 1 TB disk has approx. 268M 4 KB blocks and therefore 536M tree nodes. A cache size of 10% would hold 53.6M nodes, and with 72 B nodes (Section VII-A), would occupy 3.8 GB of memory.
(3) Optimal Hash Tree Construction:
The generated workloads are run through Huffman’s algorithm offline to produce an optimal hash tree. The tree is then serialized to disk to prepare to run the traces against it.
(4) Trace Replay:
A trace is then loaded into memory and replayed against each hash tree construction in a serial fashion. System caches are flushed before each experiment to reduce interference between experiments.
(5) Analysis:
Examining the performance of each hash tree design across a design space parameterized by capacity, cache size ratio, read/write ratio, and workload skewness allows us to understand where certain design points perform well or break down.
We defer analysis with the optimal tree oracle to Section VII.
VI. Dynamic Hash Trees
State-of-the-art hash tree designs rely on static, balanced tree structures (Android, 2023; Taassori et al., 2018), and balanced trees are specifically optimized for uniform access patterns. However, real-world storage workloads most often exhibit skewed access patterns (Li et al., 2023; Wang et al., 2023; Cooper et al., 2010; Yang and Zhu, 2016; Arasu et al., 2021). Consequently, state-of-the-art designs leave a significant amount of performance on the table (Section VII). We hypothesize that we can use reference locality as a lever to achieve higher performance. This section builds on our observations of optimal trees to explore a novel hash tree design that exploits reference locality by learning and adapting to workload patterns on-the-fly.
@@ VI-A. Challenges
Finding a Suitable Tree Structure. Unlike the balanced hash trees used in state-of-the-art designs, optimal hash trees produced by Huffman codes are often far from balanced. Figure 5 shows an example under a Zipf(2.5) probability distribution over 8192 blocks (32 MB disk). In a balanced tree, leaf node heights should be constant at 13 levels. We see two distinct regions representing hotter (height $\approx 10$) and colder data (height $\approx 30$).
@@
Yet, workloads characteristics can also vary over time: access patterns may still be skewed but regions of interest may change, or some periods of time may be characterized by more uniform access patterns. This is particularly true for storage that is shared by multiple cooperating applications or users. Thus, a tree that is optimal at one point in time may not be optimal for another (i.e., dynamically optimal). An online solution therefore must not only be able to capture hot data by placing more frequently accessed nodes higher in the tree, but also be able to dynamically adapt to changes in what particular data is deemed hot or cold over time.
@@
Online tree optimization solutions have been widely studied, particularly for search. Most algorithms focus on keeping trees balanced to reduce worst-case running time. We explicitly aim to remove the balance constraint; commonly used self-balancing trees such as AVL trees are therefore ill-fit for our use case. Instead, we aim for a more aggressive optimization: to allow the tree to become unbalanced as necessary, but be driven by the workload.
zipf2.5 Refer to caption
@@
Figure 5. Storage workloads are skewed (top) (Li et al., 2023), which suggests that the operations on the hash tree are also heavily skewed towards particular regions. This manifests in optimal hash trees (bottom) often being from balanced: under a Zipf(2.5) access probability distribution, the leaf heights show distinct regions of relatively hotter and colder data.
@@
Managing Restructuring Costs. While intuitively it makes sense that unbalanced trees should be able to exploit skewed patterns by placing more frequently accessed leaf hashes closer to the root, realizing this is in a real system is a non-trivial problem. The mechanics of adapting trees involve a series of rotations. While rotations are cheap for search trees, consisting of a series of pointer updates, they are expensive for hash trees, as we have to recompute hashes for all nodes from the rotation point up to the root; this applies when nodes are rotated during either verifications or updates.
@@ VI-B. Randomized Splaying
We draw a connection to a widely used data structure in garbage collection and IP routing lookup: splay trees (Sleator and Tarjan, 1985). Splay trees are a type of binary search tree that brings an accessed leaf to the root through a series of rotations. Importantly, splay trees capture temporal locality by keeping frequently accessed nodes closer to the root (Figure 6). However, naively used, the cost of splaying can be extremely expensive, and splaying too frequently or opportunistically may keep the tree more balanced than desired. We adapt the conventional splay tree design to meet the constraints of a hash tree. Our heuristic algorithm consists of three parameters and is performed in two steps.
@@
If the splay window is active, the splay probability denotes the probability that an accessed node should be splayed. The key intuition is that splaying is an expensive operation, but certain workload characteristics may be discernible by only examining and splaying nodes on a random subset of accesses. Finally, if a node is decided to be splayed, the splay distance defines the maximum number of levels that the node should be splayed (i.e., promoted) up the tree. This entire process occurs at the end of each verification or update call and before anything is returned to the caller. We note that this task could be offloaded to a background thread, but would trade-off giving the caller a timely integrity check, which may violate integrity guarantees; we defer closer examination of this issue to future work.
@@ VI-C. Technical Approach
We define the hotness of any tree node as the relative access frequency of the node with respect to other nodes. We track hotness of both leaves (i.e., blocks) and internal nodes (which indicate the hotness of particular subtrees/blocks). We attach an integer hotness counter to each tree node that is incremented whenever a node is promoted in the tree and decremented whenever a node is demoted in the tree. Thus, the hotness counter serves as a proxy that not only captures the relative access frequencies of nodes, but also the relative height differences in the tree as nodes are rotated.
The counter is initialized to zero after the node is authenticated and cached; the hotness of nodes that are not currently cached in memory is therefore not tracked. This allows us to localize our analysis of data hotness to the working set. Note that this approach can significantly affect performance for small caches, as it will be difficult to draw a contrast between relatively hotter, warmer, and colder data when counters are reset frequently. Nonetheless, the splay distance is a function of the hotness. The splay distance is computed in a straightforward manner: at a distance proportional to the hotness. For simplicity, we set the splay distance to be $h$ levels, where $h$ is the current hotness value of the accessed leaf.
Intuitively, so long as the hotness counters reflect the relative access frequencies of nodes, nodes accessed more frequently will be splayed further towards the root. Moreover, note that our initial exploration into this space could be expanded with sketching algorithms, machine learning, or other more sophisticated techniques (Hashemi et al., 2018).
Promotion & Demotion. After computing the splay distance, the final step is to execute the splay operation. Splaying a DHT is done in nearly the same way as it is in a search tree. There are three cases to consider when splaying a node: zig, zig-zig, and zig-zag (Figure 6). In a zig case, the node’s parent is the root, and we rotate the node up to the root. In a zig-zig case, the node’s parent is not the root, and the node and the node’s parent are either both left or right children. In this case, we perform two rotations along the same direction to rotate the node up two levels. In the zig-zag case, the node’s parent is not the root, and the node and the node’s parent are opposite-side children. Thus, we perform two rotations along opposite directions to rotate the node up two levels.
@@
Finally, splaying naturally introduces inconsistency into the tree, as it alters parent-child relationships: altering the tree structure will cause any subsequent hash fetches on a cache miss to fail due to an inconsistent root hash. We must therefore ensure that the tree remains in a consistent state by preemptively fetching (and authenticating) all auxiliary hashes before performing a rotation, then using them to commit the change immediately after. That is, parent hashes up to the root are recomputed per rotation. While updates can be costly, splaying can reduce these costs over time.
@@ VII. Evaluation
Below we evaluate the performance of DHTs in comparison to two insecure baselines (encryption/no integrity, no encryption/no integrity) and two state-of-the-art hash tree designs (binary trees used by dm-verity (Android, 2023; Chakraborti et al., 2017) and high-degree trees used by VAULT (Taassori et al., 2018)). We also discuss performance in the context of the optimal hash tree introduced in Section IV. An overview of our parameters is shown in Table I.
@@ VII-A. Experiment Setup
Read/write ratio Distribution of reads and writes
@@
TABLE I. We evaluate scalability by varying key parameters.
@@
Implementation. We implement all three hash tree designs in 4 K lines of C++. We use the BDUS framework to implement custom virtual block devices (Faria et al., 2021). BDUS provides a virtual block device abstraction layer for file systems or application software on standard Linux systems. It has a kernel module that exposes block layer hooks to userspace, and a userspace library that exposes a simple API to application code. The two primary functions of interest are read() and write(), which are invoked by the kernel whenever a block is read from or written to the block device, respectively. We perform a verify operation immediately after a block is read and an update operation immediately before a block is written to disk. Note that we consider a hash tree at the block layer like many prior works (Khati et al., 2017; Brož et al., 2018; Chakraborti et al., 2017; NIST, 2023); our basic data unit therefore aligns with the disk I/O size (4 KB).
@@
Testbed. We perform all experiments in a real cloud setting using AWS EC2 m6id.8xlarge instances, which are equipped with 32 logical cores, 128 GB memory, and 2 TB locally-attached NVMe SSDs operating at full bandwidth. We provision additional AWS EBS SSD volumes as necessary to match the capacity specified per experiment. We reinitialize hash trees between each experiment, use a standard LRU cache replacement policy, and we set the splay window flag $w=True$ and splay probability $p=0.01$ for DHTs.
@@
Cryptographic Settings. While our focus is on integrity protection, our broader goal is trusted cloud storage, and we therefore also encrypt block data in addition to authenticating it. While there are many possible cipher modes and settings (Khati et al., 2017; Brož et al., 2018), like similar works (Avanzi et al., 2022; che Tsai et al., 2017; Taassori et al., 2018) we ensure deterministic authenticated encryption through AES running in GCM mode. We use a 128-bit encryption key for blocks. The MACs produced during the encryption process are used as the leaves in the hash tree. For internal nodes, we compute 256-bit HMACs using SHA-256 with a 256-bit key.
Refer to captionRefer to caption
Refer to captionRefer to caption
zipf2.5 (cache=10%, rratio=1%, I/O size=4KB), write performance (top), read performance (bottom)
Refer to caption Refer to caption alibaba_4 (cache=10%), write performance
Figure 7. Comparison of DHTs, balanced trees, the optimal, and two insecure baselines across various capacities. DHTs provide up to a 4.2$\times$ speedup over state-of-the-art designs.
Workloads. Storage workloads are known to have a skewed shape (Li et al., 2023). Prior works benchmarking storage systems rely heavily on standard tool suites like fio or YCSB to generate skewed workloads, which primarily follow a Zipfian distribution (Yang and Zhu, 2016; Cooper et al., 2010; Arasu et al., 2021). Like these other works, we generate a set of Zipfian workloads with fio and record the traces such that we can compute an optimal tree and replay them against each hash tree design. We note that optimal trees and DHTs see qualitatively similar results across different Zipfian parameters, with larger speedups observed under increased skewness. For Zipfian workloads, we experiment with a parameter of 2.5, but note that prior measurement studies have characterized workloads as having Zipfian shape with parameters ranging from 1 to 5 (Yang et al., 2021). We note that DHTs and balanced hash trees observe similar performance under uniform workloads and focus our analysis on non-uniform workloads, which are most often seen in practice.
We also supplement analysis of Zipfian workloads with a recent dataset of block-level traces published by Alibaba in 2023 (Li et al., 2023). The traces are recorded over a set of 1000 logical volumes with different capacities, and each record in a trace is a tuple of the form $(op,offset,length)$. We perform a pre-processing step to prepare the traces for use in our evaluation framework, which requires scaling the I/O offsets and lengths to the capacity specified per experiment. For our experiments, we randomly sample one device (logical volume ID 4) from the dataset, but note that the characteristics of each volume are qualitatively the same: the mean write ratio is ¿98% (due to higher-level caches likely handling reads, and only write requests causing I/O) and the traces are heavily skewed towards small regions of the disk. Note that we focus analysis of these traces on write performance, because there are too few read requests to draw any meaningful conclusions.
State-of-the-art & Insecure Baselines. We focus our analysis of the state-of-the-art and DHTs against the optimal and two insecure baselines (no encryption/integrity, no encryption/no integrity). We note that our focus is on storage, and the state-of-the-art for storage systems is dm-verity-style trees—static, balanced, binary trees. The state-of-the-art for secure memories relies on static, balanced, high-degree (e.g., 64-ary) trees (Taassori et al., 2018). The intuition is that a larger node degree will reduce the tree height and thus the total hashing costs. Owing to their success in memory systems, we also examine high-degree trees for storage for juxtaposition.

### v3

@@ V-A. Optimal Definition
We observe that the problem of finding an optimal hash tree can be reduced to finding an optimal prefix tree (or prefix code) in the context of lossless data compression (moffat2019huffman, ). Prefix codes map a set of symbols onto a set of codewords, with the goal of compression being that codewords are as short as possible to produce a maximally compressed representation of the original data. An example is shown in Figure 7. Formally:
@@
A hash tree constructed as an optimal prefix code is optimal for an i.i.d. access probability distribution.
@@
Let $A=\{a_{1},a_{2},\ldots,a_{n}\}$ be a symbol alphabet and $W=\{w_{1},w_{2},\ldots,w_{n}\}$ be a set of associated symbol weights. Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be a prefix code that represents the set of codewords for symbols in $A$. A prefix code $C$ is said to be optimal if it minimizes the expected codeword length: $\operatorname*{arg\,min\,}_{C}~{}\sum_{i=1}^{n}w_{i}|c_{i}|,c_{i}\in C$. The length of a codeword is the number of bits in a codeword, or equivalently, the number of edges in the path from the root to the symbol leaf in the prefix tree representation of $C$. Huffman coding is a widely-used algorithm to produce optimal prefix codes (huffman1952method, ).
@@
V-B. Extended Optimal Definition
Now we extend our optimal definition to consider the effects of cache performance. Note that both data blocks and hashes can be cached in memory. In the compression domain, codeword paths in an (optimal) prefix tree can be parsed in constant work per edge, giving a total amount of work: $\sum_{i=1}^{n}w_{i}(|c_{i}|\cdot O(1))=O(1)\cdot\sum_{i=1}^{n}w_{i}|c_{i}|$. However computing a hash requires (at least) two hash fetches in a binary hash tree: the node’s two children. If both hashes are present in memory, fetch costs are negligible and the amount of work is similarly optimal: $\sum_{i=1}^{n}f_{i}(|b_{i}|\cdot O(1))=O(1)\cdot\sum_{i=1}^{n}f_{i}|b_{i}|$. Otherwise if they must be fetched from disk, I/O costs are non-negligible: $\sum_{i=1}^{n}f_{i}(|b_{i}|\cdot t(b_{i}))=\sum_{i=1}^{n}f_{i}|b_{i}|\cdot t(b_{i})$, for some function $t(b_{i})$.
@@ V-B. Extended Optimal Definition
We also see that with an optimal cache ($m=0.0$), the expected total amount of work is exactly optimal. However, it has been empirically observed that as cache size decreases, miss rates increase with a power law (chow1974optimization, ), and thus as the cache size decreases, expected I/O costs increase with a power law. This implies that the performance of hash trees is very sensitive to cache size. In particular, as cache memory can be financially costly on cloud servers, being able to synergize well with relatively smaller caches (w.r.t. larger disks) is critical to a practical hash tree deployment.
@@
Moreover, a Huffman tree is optimal under a known and fixed set of weights (cf. access probability distribution) for an i.i.d. source. If the symbol sequence (cf. block access sequence) observed while compressing a message (cf. during a program trace) exactly matches the one used to construct the tree, then the tree will be exactly optimal (i.e., provide optimal throughput). However, if the sequence deviates from the one used to construct the tree, the tree will not be exactly optimal. Similarly, if the source is not i.i.d., then the tree will not be exactly optimal—temporal patterns in the workload may cause the tree to underestimate an upper bound.
V-C. Optimal Tree Oracle
@@
Our optimal definition provides that, if we have knowledge of a concrete block access sequence (i.e., workload trace), we can instantiate an optimal hash tree from the trace and measure a concrete upper bound on performance (i.e., maximum possible throughput under the given workload). In an offline setting, where we have access to workload traces (e.g., recorded with tools like blktrace or fio), we can feasibly do so. We refer to this methodology as the optimal tree oracle.
@@
The primary purpose is to measure whether overheads observed by a hash tree design (like dm-verity) are better attributed to the structure of the tree or to a fundamental scaling limit. For example, a hash tree may be performing optimally, but have high overheads, which would suggest that complimentary optimizations (e.g., dividing the tree into one or more independent security domains) may be the only way to break the performance ceiling. In contrast, a hash tree that does not perform optimally under a given workload may require a fundamental redesign.
@@
We liken this approach to Belady’s optimal page replacement algorithm (belady1966study, ), a clairvoyant algorithm that has a priori knowledge of future memory accesses and can make optimal page replacement decisions. This gives us the ability to make rigorously grounded conclusions about what hash tree designs perform well and when. We defer analysis with the optimal tree oracle (denoted by H-OPT) to Section VII.
VI. Dynamic Merkle Trees
A condition of instantiating an optimal hash tree is that we must have a priori knowledge of the exact workload. This is rarely feasible in practice. This section builds on observations of optimal hash trees to develop a novel hash tree design that can approximate an optimal tree by learning and adapting to workload patterns on-the-fly.
zipf2.5 Refer to caption
Figure 8. In practice, access patterns observed at the storage layer are most often skewed, and Zipfian workloads are used to model this shape (li2023depth, ; yang2016write, ). This graph shows that a small number of blocks is accessed most of the time, which suggests that operations on the hash tree should also be skewed.
@@ VI-A. Challenges
Finding a Suitable Tree Structure. State-of-the-art hash tree designs rely on static, balanced tree structures (android-dm-verity, ; taassori2018vault, ). Balanced trees are optimal under uniform access patterns. However, real-world storage workloads most often exhibit skewed (i.e., non-uniform) access patterns (li2023depth, ; wang2023locality, ; cooper2010benchmarking, ; yang2016write, ; arasu2021fastver, ) where a small number of blocks is accessed most of the time. The result is that the optimal hash trees produced by Huffman codes are often far from balanced.
For example, Figure 8 shows the access distribution for a Zipfian workload; note that most real-world workloads obey Zipf’s law and are modeled with Zipfian workloads (yang2016write, ). The data accesses are highly skewed, which suggests that the operations on the hash tree may also be highly skewed. Figure 9 shows this to be true: in a balanced tree over 8192 blocks (a 32 MB disk), leaf node heights are constant at 13, but in the optimal tree we see two distinct regions representing hotter (height $\approx 10$) and colder data (height $\approx 30$).
Optimal trees tend to accumulate hot data high up in the tree and place cold data at nearly a 3×3$\times$3 × height difference—significantly reducing the verify/update latency for hot data. This indicates that an optimal tree is one that aggressively optimizes for hot data (the working set). Unfortunately, the static nature of standard hash tree designs precludes exploiting this skew when present in a workload.
@@
Handling Changing Access Patterns. Yet, workload characteristics can also vary over time: access patterns may still be skewed but regions of interest may change, or some periods of time may be characterized by more uniform access patterns. This is particularly true for storage that is shared by multiple cooperating applications or users. Thus, a tree that is optimal at one point in time may not be optimal for another (i.e., dynamically optimal). An online solution, one that does not assume a priori knowledge of workload characteristics, therefore must not only be able to capture hot data by placing more frequently accessed nodes higher in the tree, but also be able to dynamically adapt to changes in what particular data is deemed hot or cold over time.
@@
Adaptive tree structures have been widely studied, particularly for search. Most algorithms focus on keeping trees balanced to reduce worst-case running time. We explicitly aim to remove this constraint; commonly used self-balancing trees (e.g., AVL trees) are therefore ill-fit for our use case. We aim for a more aggressive optimization: allow the tree to become unbalanced as necessary, but be driven by the workload.
@@
Figure 9. Skewed patterns manifest in optimal hash trees being far from balanced: under a Zipfian probability distribution, there are distinct regions of relatively hotter and colder data.
@@
Managing Restructuring Costs. While intuitively it makes sense that unbalanced trees should be able to exploit skewed patterns by placing more frequently accessed leaf hashes closer to the root, realizing this is in a real system is a non-trivial problem. The mechanics of adapting trees involve a series of rotations. While rotations are cheap for search trees, consisting of a series of pointer updates, they are expensive for hash trees, as we have to recompute hashes for all nodes from the rotation point up to the root. This applies when nodes are rotated during either verifications or updates.
@@ VI-B. Randomized Splaying
We draw a connection to a data structure widely used in garbage collection and IP routing: splay trees (sleator1985self, ). Splay trees are a type of binary search tree that brings an accessed leaf to the root through a series of rotations. Importantly, splay trees capture temporal locality by keeping frequently accessed nodes closer to the root (Figure 10). However, naively used, the cost of splaying can be extremely expensive, and splaying too frequently or opportunistically may keep the tree more balanced than desired. We adapt the conventional splay tree design to meet the constraints of a hash tree.
@@
If the splay window is active, the splay probability denotes the probability that an accessed node should be splayed. The key intuition is that splaying is an expensive operation, but we can amortize costs by only splaying on a small percentage of accesses (e.g., 1% of the time). Finally, if a node is decided to be splayed, the splay distance defines the maximum number of levels that the node should be splayed (i.e., promoted) up the tree. This entire process occurs at the end of each verification or update call and before anything is returned to the caller.
@@ VI-C. Technical Approach
We attach an integer hotness counter to each tree node that is incremented whenever a node is promoted in the tree and decremented whenever a node is demoted in the tree. This applies to both leaves (i.e., blocks) and internal nodes (which indicate the hotness of particular subtrees/blocks). The purpose of the hotness counter is to track the relative access frequencies of nodes as they are rotated, and use this information to determine how far to splay the node up the tree.
The counter is initialized to zero after the node is authenticated and cached; the hotness of nodes that are not currently cached in memory is therefore not tracked. The purpose of this is to localize our analysis of data hotness to the working set. Note that this approach can negatively affect performance for small caches, as it will be difficult to draw a contrast between relatively hotter, warmer, and colder data when counters are reset frequently. Nonetheless, the splay distance is a function of the hotness. The splay distance is computed in a straightforward manner: at a distance proportional to the hotness. For simplicity, we set the splay distance to be $h$ levels, where $h$ is the current hotness counter value of the accessed leaf.
Intuitively, nodes that are deeper in the tree (colder) will climb the tree slowly, while nodes that are higher in the tree (hotter) will climb quickly. Note that our initial exploration into this space could be expanded with sketching algorithms, machine learning, or other sophisticated techniques (hashemi2018learning, ).
Promotion & Demotion. After computing the splay distance, the final step is to execute the splay operation. Splaying a DMT is done in nearly the same way as it is in a search tree. There are three cases to consider when splaying a node: zig, zig-zig, and zig-zag (Figure 10). In a zig case, the node’s parent is the root, and we rotate the node up to the root. In a zig-zig case, the node’s parent is not the root, and the node and the node’s parent are either both left or right children. In this case, we perform two rotations along the same direction to rotate the node up two levels. In the zig-zag case, the node’s parent is not the root, and the node and the node’s parent are opposite-side children. Thus, we perform two rotations along opposite directions to rotate the node up two levels.
@@
Finally, splaying naturally introduces inconsistency into the tree, as it alters parent-child relationships. This will cause any subsequent hash fetches on a cache miss to fail due to an inconsistent root hash. We must therefore ensure that the tree remains in a consistent state by preemptively fetching (and authenticating) all sibling hashes before performing a rotation, then using them to commit the change immediately after. That is, parent hashes up to the root are recomputed per rotation (see ”Update from” in Figure 10). While updates can be costly, splaying can reduce these costs over time.
@@ VII. Evaluation
We compare DMT performance against two insecure baselines (No encryption/no integrity, Encryption/no integrity), the binary trees used by dm-verity (android-dm-verity, ), and the optimal binary tree (H-OPT). We also juxtapose DMTs against the high-degree trees that have been widely used in secure memory systems (taassori2018vault, ; feng2021scalable, ). Finally, as noted previously (see Figure 6), we also examine DMTs with respect to 4-ary and 8-ary trees, which have not been considered by prior work. An overview of our parameters is shown in Table I.
@@ VII-A. Experiment Setup
Read ratio % of read operations
I/O size Size of application I/O
I/O depth Max no. outstanding application I/Os
Thread count Number of application threads
TABLE I. Experiment parameters.
@@
Implementation. We implement the hash trees in 5 K lines of C++. We use BDUS to implement custom block device drivers that wrap lower-level drivers (faria_bdus_2021, ). BDUS has a kernel module that exposes block layer hooks to userspace. The two primary functions of interest are read() and write(), which are invoked by the kernel whenever a block is read from or written to the block device, respectively. We perform a verify immediately after a block is read and an update immediately before a block is written to disk. Our basic data unit aligns with the disk I/O size (4 KB blocks) (khati2017full, ; brovz2018practical, ; chakraborti2017dm, ; nist-aes-xts, ).
@@
Testbed. We perform all experiments using AWS EC2 i4i.8xlarge instances equipped with 32 cores, 256 GB memory and locally-attached NVMe SSDs for data and metadata. Note that currently there are no available cloud instance types which both have local NVMe storage and also support confidential VM technology such as AMD SEV-SNP. We reinitialize hash trees between each experiment, use a standard LRU cache replacement policy, and we set the splay window flag $w=True$ and splay probability $p=0.01$ for DMTs.
@@
Cryptographic Settings. Like prior works, we ensure deterministic authenticated encryption with AES-GCM (avanzi2022cryptographic, ; taassori2018vault, ). We use a 128-bit encryption key for blocks. The MACs produced during the encryption process are used as the leaves in the hash tree. For internal nodes, we compute 256-bit hashes using SHA-256 with a 256-bit key.
Workload Settings. We perform a broad analysis across different system and workload configurations, parameterized by disk capacity, hash cache size, read/write ratio, I/O size, thread count, and I/O depth. This enables exhaustively examining the performance space of DMTs. We also examine different degrees of workload skewness, from pure uniform to highly skewed. We focus especially on the Zipfian workload discussed in Section IV, which closely approximates real-world block-level access patterns, which are highly skewed (and write-heavy) (li2023depth, ; yang2016write, ); see Figure 18. We also use a recently published Alibaba dataset recorded from an array of 1000 volumes backing various virtual machines in a public cloud datacenter (li2023depth, ). Finally, we demonstrate how driver-level improvements translate to application-level improvements with a case study of the Filebench OLTP workload (Tarasov2016FilebenchAF, ).
Like prior works, we generate workloads with fio (fio, ); we record/replay traces for the optimal. Workloads have a 5 minute warmup period and 15 minute benchmark period.